OpenStudy (anonymous):

The depth (inches) of water that accumulates in the spring soil from melted snow can be determined by dividing the cumulative winter snowfall (inches) by 12. Determine the amount of winter snow that accumulates 6 inches of water in the soil. a. 6 inches of snow b. 84 inches of snow c. 72 inches of snow d. 12 inches of snow

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

\[12*6=72/12=6\] this should help im not exactly positive though

OpenStudy (anonymous):

did it work?